please help i need this answer soon if u can thanks Find r if, I = $1,680, p = $12,000, t = 2 years.


Answer by mananth(16946)

Find r if, I = $1,680, p = $12,000, t = 2 years.
I=p*r*t
1680=12000*r*2
r=1680/12000*2
r=0.07 = 7%